EPLAIN(1) EPLAIN(1) NAME eplain−extendedplainTeX SYNOPSIS eplain[firstline] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. eplainisanextensionofplainTEXdescribedinthebookTEXfortheImpatient.Itaddsthese features(amongothers)toplainTEX: •left-justifieddisplays •double-columnoutput •tablesofcontents •\hruleand\vrulewithadifferentdefaultthan0.4pt •timeofday •verbatimfilelisting •generalizedfootnotes •blankandblackboxes •citationsusingBIBTEX FILES eplain/doc/eplain.texi TEXabledocumentationofetexmacrosintexinfoformat. eplain.info* OnlineGNUEmacsinfo(1)filesproducedfromeplain.texi. ThecompleteeplaindistributionisavailableviaanonymousftptotheInternethost ftp.cs.umb.edufromthefilepub/tex/eplain.tar.Zwhichislinkedtoanotherfilecontainingthe mostrecentversionofthepackage.MajorTEXarchivesitesshouldalsohaveit. SEEALSO emacs(1),info(1),initex(1),tex(1),xinfo(1). KarlBerry,eplain,TUGboat11(4),571--572,1990. PaulW.AbrahamswithKarlBerryandKathrynA.Hargreaves,TEXfortheImpatient,Addison-Wesley,1990,ISBN0-201-51375-7. Web2C7.32x 14December1993 1